Required information (The following information applies to the questions displayed below.) Summary information from the finanComplete this question by entering your answers in the tabs below. 1A Current Ratio 1A Acid Test Ratio 1A Acct Rec Turn 1A InComplete this question by entering your answers in the tabs below. 1A Current Ratio 1A Acid Test Ratio 1A Acct Rec Turn 1A InComplete this question by entering your answers in the tabs below. UUUUUUUUUUUUUUUUUUUUUUUU 1A Current Ratio 1A Acid Test RatComplete this question by entering your answers in the tabs below. 1A Current Ratio 1A Acid Test Ratio 1A Acct Rec Turn 1A InRequired: 1a. For both companies compute the (a) current ratio, (b) acid-test ratio, (C) accounts receivable turnover, (d) inRequired: 1a. For both companies compute the (a) current ratio, (b) acid-test ratio, (c) accounts receivable turnover, (d) inRequired: 1a. For both companies compute the (a) current ratio, (b) acid-test ratio, (c) accounts receivable turnover, (d) in
Answer #1

1. A) Current Ratio

Company Choose Numerator / Choose Denominator = Current Ratio
Current Asset / Current Liabilities = Current Ratio
Barco $145,940 / $70,340 = 2.07
Kyan $228,550 / $101,300 = 2.56

Current Asset Barco = $22,000+$33,400+$84,640+$5,900 = $145,940

Kyan = $35,000+$51,400+$134,500+$7,650=$228,550

B. Acid Test Ratio

Company Choose Numerator / Choose Denominator = Acid Test Ratio
Cash Accounts Receivable / Current Liabilities = Acid Test Ratio
Barco $22,000 $33,400 / $70,340 = .79
Kyan $35,000 $51,400 / $101,300 = .85

C. Accounts Receivable Turnover

Company Choose Numerator / Choose Denominator = Accounts Receivable Turnover
Net Sales / Average Accounts Receivable = Accounts Receivable Turnover
Barco $770,000 / $29,600 = 26.01
Kyan $912,200 / $52,800 = 17.28

D. Inventory Turnover

Company Choose Numarator / Choose Denominator = Inventory Turnover
Cost of Goods Sold / Average Inventory = Inventory Turnover
Barco $590,100 / $73,120 = 8.07
Kyan $638,500 / $119,950 = 5.32

E. Days Sales in Inventory

Company Choose Numarator / Choose Denominator × Days = Days Sales in Inventory
Merchandise Inventory / Cost of Goods Sold × 365 = Days Sales in Inventory
Barco $84,640 / $590,100 × 365 = 52.35 Days
Kyan $134,500 / $638,500 × 365 = 76.88Days

F. Days Sales Uncollected

Company Choose Numarator / Choose Denominator × Days = Days Sales Uncollected
Current Receivable / Net Sales × 365 = Days Sales Uncollected
Barco $33,400 / $770,000 × 365 = 15.8 Days
Kyan $51,400 / $912,200 × 365 = 20.6 Days

2. Barco Company is the Better short term credit risk